OpenStudy (anonymous):

1+3+5+7+9+11+1+15+17+19=100 then out of this which 5 numbers make 50 .some logic working?

Nnesha (nnesha):

first above series are not equal to 100 1+3+5+7+9+11+1+15+17+19=88

Nnesha (nnesha):

and after 11 there suppose to be 13 not one because there are odd numbers and after adding 13 instead of 1 then its make = to 100

Nnesha (nnesha):

and 50 is even number we cannot get 50 out of 5 odd numbers but it can be possible if we use 6 digits and also if we use another even number like 4 digits 19+13+7+11 = 50 17+15+7+11 = 50

OpenStudy (anonymous):

That is not \(1\), that is \(13\) infact.. :)

OpenStudy (anonymous):

\(1+3+5+7+9+11+13+15+17+19=100\)

OpenStudy (anonymous):

Add any five numbers and make \(50\).

Nnesha (nnesha):

100-19-17-13-1=50 so it could be well in the above question nowhere he said add " then out of this which 5 numbers make 50"